Write the state of matter for every substance in the chemical reaction.

Explain your limiting reactant determinations.

Mark every measured number showing the significant figures (SF).

Do not round any intermediate values. (you can show two digits past the marked SF)

Show subtraction and addition in the vertical format.

Show your final answer before rounding and then in the final form, rounded correctly.

Every number must have units and substances. You can use tables or other means to increase your efficiency, but the reader must be able to clearly identify the units and substances related to every number.

Include descriptive statements indicating what you are doing. For example: "Calculate the moles of the reactants"

EXERCISE 3.6.2: LANTHANUM OXALATE

Show the steps to answer what mass of solid lanthanum(III) oxalate nonahydrate [La2(C2O4)3•9H2O] can be obtained from 650 mL of a 0.0170 M aqueous solution of LaCl3 by adding a stoichiometric amount of sodium oxalate?

Strategy:

  1. Check the chemical equation to make sure it is balanced as written; balance if necessary. Then calculate the number of moles of [Au(CN)2] present by multiplying the volume of the solution by its concentration.
  2. From the balanced chemical equation, use a mole ratio to calculate the number of moles of gold that can be obtained from the reaction. To calculate the mass of gold recovered, multiply the number of moles of gold by its molar mass.

Answer

3.89 g
Answer #1

Given molarity of LaL3 solution, C = 0.0170 mol/L

Volume of  LaL3 solution taken, V = 650 mL * (1L / 1000 mL) = 0.650 L

=> Moles of LaL3 initially taken = C*V = 0.0170 mol/L * 0.650 L = 0.01105 mol

The balanced equation for reaction of  LaCl3 and sodium oxalate, Na2C2O4 to form solid lanthanum(III) oxalate nonahydrate, La2(C2O4)3.9H2O(s) is:

2LaCl3(aq) + 3Na2C2O4(aq) +9H2O(l) ----> La2(C2O4)3.9H2O(s) + 6NaCl(aq)

The coefficient of LaCl3(aq) in the balanced equation is 2 and that of La2(C2O4)3.9H2O(s) is 1.

Hence mole ration between La2(C2O4)3.9H2O(s) and LaCl3(aq) is 1:2.

=> Moles of La2(C2O4)3.9H2O(s) formed = 0.01105 mol LaCl3(aq) * [1 mol La2(C2O4)3.9H2O(s) / 2 mol LaCl3(aq) ]

= 0.01105 * (1/2) mol  La2(C2O4)3.9H2O(s)

= 0.005525 mol La2(C2O4)3.9H2O(s)

Molar mass of La2(C2O4)3.9H2O(s) = 704.0 g/mol

=> Mass of La2(C2O4)3.9H2O(s) obtained = 0.005525 mol * (704.0 g / 1 mol) = 3.89 g (Answer)